# FedLPA: One-shot Federated Learning with Layer-Wise Posterior Aggregation

### v3

@@ Appendix I Expectation approximation (EA)
Following the differential privacy (DP) mechanisms [77, 86, 78, 87] to protect privacy, we conduct the two mechanisms of DP-FedLPA: (1) adding Laplace random noise to the training data samples, (2) adding Laplace random noise to the parameters to be transmitted. DP is a rigorous and popular privacy metric, which guarantees that the output does not change with a high probability even though an input data record changes. Specifically, since the sensitivity of the data sample distribution after the normalization is 1, we add Laplacian noises with $\lambda=\frac{1}{\epsilon}$. We set $\epsilon=\{3,5,8\}$ that provides modest privacy guarantees since normally $\epsilon\in(1,10)$ is viewed as a suitable choice. We have added the experiments using the same experiment setting in the paper with three random seeds and 10 clients on the FMNIST dataset. Results are shown in Table 7. DP-FedLPA under both mechanisms outperforms FedAvg, which shows that it is compatible with combining our proposed FedLPA with DP to enhance privacy protection levels. Note that the smaller $\epsilon$ is, the larger noises we add. We find that when the $\epsilon$ gets smaller, the performance drops simultaneously, while the privacy protection level is increased.
@@
8 $\beta$=0.1 30.93$\pm$0.58 50.01$\pm$0.07 57.15$\pm$1.23
$\beta$=0.3 45.27$\pm$0.05 68.30$\pm$0.41 66.21$\pm$0.14
$\beta$=0.5 59.10$\pm$0.63 71.17$\pm$0.27 73.50$\pm$0.06
5 $\beta$=0.1 30.93$\pm$0.16 48.51$\pm$0.07 55.87$\pm$ 0.88
$\beta$=0.3 45.27$\pm$0.05 67.34$\pm$0.92 66.02$\pm$0.71
$\beta$=0.5 59.10$\pm$0.63 70.89$\pm$0.80 73.44$\pm$0.20
3 $\beta$=0.1 30.93$\pm$0.79 48.39$\pm$0.07 54.31$\pm$0.44
$\beta$=0.3 45.27$\pm$0.05 65.08$\pm$0.45 65.22$\pm$0.46
$\beta$=0.5 59.10$\pm$0.63 70.28$\pm$1.30 72.19$\pm$0.62
L.2 Experiments with iDLG attack
@@
L.3 Concrete examples of privacy attack
@@
Appendix M Additional experiments
M.1 Dataset information
Our experiments were conducted using four real-world datasets: MNIST [55], FMNIST [56], SVHN [58], and CIFAR-10 [57]. Here’s a brief overview of these datasets:
MNIST Dataset: The MNIST dataset comprises binary images of handwritten digits. It consists of 60,000 28x28 training images and 10,000 testing images.
FMNIST Dataset: Similar to MNIST, the FMNIST dataset also contains 60,000 28x28 training images and 10,000 testing images.
SVHN Dataset: The SVHN dataset includes 73,257 32x32 color training images and 10,000 testing images.
CIFAR-10 Dataset: CIFAR-10 consists of 60,000 32x32 color images distributed across ten classes, with each class containing 6,000 images.
The input dimensions for MNIST, FMNIST, SVHN, and CIFAR-10 are 784, 784, 3,072, and 3,072, respectively.
M.2 t-SNE visualization
@@
M.3 Experiments on different local epoch numbers
@@
M.4 Extreme setting, 5 clients
@@
M.5 Extreme setting, $\beta=0.001$
@@
M.6 Aggregation visualization
Figure 6: Aggregation Visualization.
Figure 6 displays the visualization aggregation of two heterogeneous Gaussians. In comparison to the Fisher Diagonal [17] method and FedAvg, our method aggregates the local clients on the Riemannian manifold while treating the empirical Fisher information matrix as a metric of the parameter space, which is much better.
M.7 Experiments with FedOV and Co-Boosting
@@
In conclusion, our method could be comparable with FedOV and Co-Boosting in some settings, even when they consume more computational resources as shown in Appendix M.9.
@@
M.8 Communication overhead evaluation
@@
M.9 Running time and computation overhead evaluation
@@
M.10 Experiments with more complex neural network structure
We do the experiments with FedLPA on the same experiment setting in the paper using the more complex network, ResNet-18 [97], with three random seeds on the CIFAR-10 dataset. We set the parameters with $\beta$=0.1, 0.3 and 0.5 with 10 clients. The results are shown in Table 18. From the results, we could see that using ResNet-18, our method still gets better performance compared to other baselines.
We do the experiments with FedLPA on the same experiment setting in the paper using the more complex network, VGG-9 [98], with three random seeds on the FMNIST dataset. We set the parameters with $\beta$=0.1, 0.3 and 0.5 with 10 clients. The results are shown in Table 19. From the results, we could see that using VGG-9, our method still performs better than other baselines.
@@
M.11 Experiments with more complex datasets
We did the experiments with FedLPA on the same experiment setting in the paper using ResNet-18 with three random seeds on the CIFAR-100 [57] dataset. The results are shown in Table 20. We can see that even with the complicated dataset CIFAR-100, our method could also get satisfactory results in the federated one-shot setting. Besides, we also have added the experiments on EMNST [99] using simple-CNN with 10 clients and three random seeds. We do the experiments on EMNIST-mnist and EMNIST-letters. The results are shown in Table 21.
@@
M.12 Ablation experiments analyzing the number of approximation iterations of FedLPA
@@
(2) Block-diagonal Fisher matrix to approximate the full one:
The work [101] provides a detailed evaluation and testing of using block-diagonal Fisher to approximate the full one. Firstly, Chapter 6.3.1 “Interpretations of this approximation" in the paper [101] indicates that using a block-wise Kronecker-factored Fisher closely approximates the full Fisher. Although there is a bias term (due to the approximation in Eq. 30 ), this term approximates zero when there are sufficient samples. Furthermore, the paper examines the approximation quality of block-diagonal Fisher compared with the true Fisher and suggests that block-diagonal Fisher captures the main correlations, while the remaining correlations have a minimal impact on the experimental results.
(3) Besides, we have added some experiments for more ablation studies with our method on the same experiment setting in the paper with three random seeds with 10 clients. We conducted experiments on FMNIST datasets with $\beta$=0.1, 0.3 and 0.5. The results are shown in Table 22. We show the experiments analyzing the number of approximation iterations. With the experiment results, we could know that 5000 iterations are enough to get the ideal results. By default, we use 10000 iterations.
@@
In Appendix J, we also show the linear convergence rate for Eq. 14.
@@
M.13 Experiments with different proportion of data samples
We have added the experiments with our method on the same experiment setting in the paper with three random seeds with 10 clients. We conducted experiments on FMNIST datasets with $\beta$=0.1, 0.3 and 0.5. The performance change w.r.t the number of data samples are shown in Table 23. We could see that our method FedLPA could yield satisfactory results even with only 20% data samples under multiple settings.
Table 23: Experiments with different proportions of data samples.
100% 55.33$\pm$0.06 68.20$\pm$0.04 73.33$\pm$0.06
80% 53.88$\pm$1.14 65.47$\pm$0.02 73.17$\pm$0.05
60% 53.15$\pm$0.82 64.80$\pm$0.71 72.40$\pm$0.29
40% 53.20$\pm$0.21 64.10$\pm$0.40 70.02$\pm$0.17
20% 45.71$\pm$0.13 62.15$\pm$0.03 68.54$\pm$2.02
M.14 Artifact details
@@
The experiments_resnet.sh script covers the experiments in Table 18 and Table 20 and takes about 25 days.
The experiments_vgg.sh script covers the experiments in Table 19 and takes about 7 days.

### v5

@@ F.1 Experiments with differential privacy
Following the differential privacy (DP) mechanisms [56, 85, 57, 86] to protect privacy, we conduct the two mechanisms of DP-FedLPA: (1) adding Laplace random noise to the training data samples, (2) adding Laplace random noise to the parameters to be transmitted. DP is a rigorous and popular privacy metric, which guarantees that the output does not change with a high probability even though an input data record changes. Specifically, since the sensitivity of the data sample distribution after the normalization is 1, we add Laplacian noises with $\lambda=\frac{1}{\epsilon}$. We set $\epsilon=\{3,5,8\}$ that provides modest privacy guarantees since normally $\epsilon\in(1,10)$ is viewed as a suitable choice. We have added the experiments using the same experiment setting in the paper with five random seeds and 10 clients on the FMNIST dataset. Results are shown in Table 8. DP-FedLPA under both mechanisms outperforms FedAvg, which shows that it is compatible with combining our proposed FedLPA with DP to enhance privacy protection levels. Note that the smaller $\epsilon$ is, the larger noises we add. We find that when the $\epsilon$ gets smaller, the performance drops simultaneously, while the privacy protection level is increased.
Besides, we have added the experiments using the same experiment setting to show the round results of how many rounds DP-FedAvg needs to achieve the same test performance with the first mechanism. The results in Table 9 show that DP-FedAvg needs about 10 rounds of communication to achieve the same test performance, compared to our one-round FedLPA. Combined with our previous results in Table 5 and Table 8, our FedLPA could save the communication and computation overhead and combine with the DP method to mitigate the potential privacy leakage. Based on the above settings, DP-FedAvg needs at least 3x communication overhead and 5x computation overhead. While DP-FedAvg needs multiple rounds to get similar accuracy, DP-FedAvg maybe vulnerable to more privacy attack methods due to the multiple queries, such as curvature-based privacy attacks.
@@
8 $\beta$=0.1 31.90$\pm$0.58 50.01$\pm$0.07 57.15$\pm$1.23
$\beta$=0.3 44.37$\pm$0.05 68.30$\pm$0.41 66.21$\pm$0.14
$\beta$=0.5 57.92$\pm$0.63 71.17$\pm$0.27 73.50$\pm$0.06
5 $\beta$=0.1 28.17$\pm$0.16 48.51$\pm$0.07 55.87$\pm$ 0.88
$\beta$=0.3 43.91$\pm$0.05 67.34$\pm$0.92 66.02$\pm$0.71
$\beta$=0.5 57.14$\pm$0.63 70.89$\pm$0.80 73.44$\pm$0.20
3 $\beta$=0.1 27.85$\pm$0.79 48.39$\pm$0.07 54.31$\pm$0.44
$\beta$=0.3 42.80$\pm$0.05 65.08$\pm$0.45 65.22$\pm$0.46
$\beta$=0.5 54.80$\pm$0.63 70.28$\pm$1.30 72.19$\pm$0.62
Table 9: Experiments with Differential Privacy for Round Numbers.
$\beta$ $\epsilon$ 8 5 3
0.1 11 11 12
0.3 11 9 8
0.5 8 8 7
F.2 Experiments with iDLG attack
@@
F.3 Concrete examples of privacy attack
@@
Appendix G Additional experiments
G.1 t-SNE visualization
@@ Appendix G Additional experiments
G.2 Experiments on different local epoch numbers
@@
G.3 Extreme setting, 5 clients
@@
G.4 Extreme setting, $\beta=0.001$
@@
G.5 Experiments with FedOV and Co-Boosting
@@
In conclusion, our method could be comparable with FedOV and Co-Boosting in some settings, even when they consume more computational resources as shown in Appendix G.7.
@@
G.6 Communication overhead evaluation
@@
G.7 Running time and computation overhead evaluation
@@
G.8 Experiments with more complex neural network structure
We do the experiments with FedLPA on the same experiment setting in the paper using the more complex network, ResNet-18 [95], with five random seeds on the CIFAR-10 dataset. We set the parameters with $\beta$=0.1, 0.3 and 0.5 with 10 clients. The results are shown in Table 20. From the results, we could see that using ResNet-18, our method still gets better performance compared to other baselines.
We do the experiments with FedLPA on the same experiment setting in the paper using the more complex network, VGG-9 [96], with five random seeds on the FMNIST dataset. We set the parameters with $\beta$=0.1, 0.3 and 0.5 with 10 clients. The results are shown in Table 21. From the results, we could see that using VGG-9, our method still performs better than other baselines.
@@ G.8 Experiments with more complex neural network structure
G.9 Experiments with more complex datasets
We do the experiments with FedLPA on the same experiment setting in the paper using ResNet-18 with five random seeds on the CIFAR-100 [65] dataset. The results are shown in Table 22. We can see that even with the complicated dataset CIFAR-100, our method could also get satisfactory results in the federated one-shot setting. Besides, we also have added the experiments on EMNST [97] using simple-CNN with 10 clients and five random seeds. We do the experiments on EMNIST-mnist and EMNIST-letters. The results are shown in Table 23.
In addition to these, we conduct experiments with Tiny-ImangeNet [98] with ResNet-18 with 10 clients and five random seeds. The results are shown in Table 24.
@@ G.9 Experiments with more complex datasets
Table 24: Experiments with Tiny-ImangeNet using ResNet-18.
$\beta$ FedLPA Dense FedAvg
0.1 17.02$\pm$1.40 15.88$\pm$1.96 3.72$\pm$1.44
0.3 27.80$\pm$2.10 24.91 $\pm$1.65 8.41$\pm$0.87
0.5 30.14$\pm$1.25 29.43$\pm$0.72 12.07 $\pm$1.92
G.10 Ablation experiments analyzing the number of approximation iterations of FedLPA
@@
(2) Block-diagonal Fisher matrix to approximate the full one: The work [100] provides a detailed evaluation and testing of using block-diagonal Fisher to approximate the full one. Firstly, Chapter 6.3.1 “Interpretations of this approximation" in the paper [100] indicates that using a block-wise Kronecker-factored Fisher closely approximates the full Fisher. Although there is a bias term, this term approximates zero when there are sufficient samples. Furthermore, the paper examines the approximation quality of block-diagonal Fisher compared with the true Fisher and suggests that block-diagonal Fisher captures the main correlations, while the remaining correlations have a minimal impact on the experimental results.
(3) Besides, we have added some experiments for more ablation studies with our method on the same experiment setting in the paper with five random seeds with 10 clients. We conducted experiments on FMNIST datasets with $\beta$=0.1, 0.3 and 0.5. The results are shown in Table 25. We show the experiments analyzing the number of approximation iterations. With the experiment results, we could know that 5000 iterations are enough to get the ideal results. By default, we use 10000 iterations.
@@
G.11 Artifact details